# Learned feature representations are biased by complexity, learning order, position, and more

### v1

@@ 5 Vision
We round out our experiments by considering a third large class of problems—computer vision—and the architectures typically used to solve them. Specifically, we construct synthetic datasets of images that show objects with distinct colors, shapes, and textures, and train ResNets (He et al., 2016) and other CNNs without residual connections to classify all of these features simultaneously. These experiments are again inspired by prior work finding that features like color are easier for models to learn, and will dominate when multiple features compete to predict a label (Hermann and Lampinen, 2020). Here, we investigate the models’ internal representations of these features when all features are learned.
We find that non-residual CNNs exhibit a strong bias towards the simplest feature (color, which can be decoded from any single pixel) at the final spatial pooling layer, and a somewhat weaker bias at the penultimate classifier layer (Fig. 11(a)). This is in keeping with the general pattern observed above, that models are biased towards representing simpler features more strongly, as well as with prior work
Intriguingly, however, ResNets exhibit the opposite bias—with representations that strongly favor shape (Fig. 11(b)). This is despite the fact that color remains the easiest feature, and shape the most difficult (slowest) feature to learn for both models (Fig. 36; as in Hermann and Lampinen 2020).
@@ 5.1 Relational vision tasks
Why do the basic feature biases of ResNets change direction in these tasks, compared to those above? In follow-up experiments (Appendix B.9.1) we show that switching to binary classification, and the addition of other objects to the image, may both contribute to reversing these biases. These results highlight the complex interactions that affect representational biases. However, the bias towards basic over relational features seems relatively more consistent across architectures and feature types.
@@ 6.1 The representations identified are causally sufficient and specific
The above analyses are correlational. However, most philosophical accounts (e.g. Cao, 2022; Baker et al., 2022) define representations partly in terms of their causal role in a computation. An increasing number of approaches in interpretability similarly use causal interventions to achieve more faithful interpretations of model internal representations (e.g. Geiger et al., 2021; Meng et al., 2022; Geiger et al., 2024). In this section, we verify in one case that the representations identified in our analyses are indeed playing a causal role in the computations.
To do so, we take an MLP model (from the Fig. 2 experiments) at the end of training, manipulate its internal representations, and evaluate the causal effect on its output behavior. To manipulate the model representations of a feature, we take the representation pattern predicted when that feature is present or absent and subtract those two patterns to create a present-absent difference vector. We then add (or subtract) a scalar multiple of this difference vector to the representation of each stimulus in the test dataset, and evaluate whether by doing so we can flip the label the model predicts for that feature on that stimulus.
@@ 6.2 Impact on model simplification for interpretability
Figure 14: If we keep only the top-$k$ principal components from a models representations, the easy feature will be preserved with a smaller $k$ than the hard feature.
@@
Though our experimental settings are somewhat different, our results hint at a possible basis for those phenomena. In particular, if the features required to solve the harder OOD splits are more complex, they may occupy relatively less variance, and therefore be more readily lost when keeping only the top principal components. We show a simple demonstration of this from one of our MLP models in Fig. 14: if we keep only two principal components from the models representations, the model can still identify the easy feature perfectly, while hard feature accuracy remains at chance until we keep at least four principal components. (In Appx. B.10 we show that dropping the top PCs yields the opposite pattern, with hard features proving more robust).
@@ 6.2.1 Impact on feature visualization
Given the above observations, it is clear that low-dimensional visualization may be biased towards certain features. We showed a simple example of this in Fig. 3, in which the top PCs are entirely clustered by the easy feature. In Appx. B.11 we elaborate these results, showing that the hard feature appears in the next principal component, but if models are trained with more and more easy features, the hard feature is pushed into higher and higher PCs. Visualizing with $t$-SNE yields somewhat better results when there is only a single easy feature—due to $t$-SNE’s nonlinear embedding and emphasis on neighbors—but once enough easy features are added it similarly becomes difficult to identify the hard feature in low-dimensional $t$-SNE plots as well, while the easy features show extremely clear clusters. Thus, feature visualizations may be inherently biased towards simpler features, those that are learned earlier, etc.
@@ 6.3 Impact on Representational Similarity Analysis
To illustrate this, in Fig. 15 we perform an RSA analysis between many pairs of MLP networks. Each MLP is trained to output two features of varying difficulty. Ideally, the RSA would be more similar on the diagonal (for networks trained to perform the same two tasks) than off the diagonal, and those on the block diagonals (trained with just one feature that is the same) would be more similar than those off. However, the feature representation biases lead to a substantially more biased structure. For example, two models that are both computing SUM%2 over two sets of features will appear to be less similar to one another than each appears to be a network that is solely computing simple linear features. These results use a Euclidean metric for RSA, which is naturally sensitive to variance. However, in Appx. B.12 we show that the results are qualitatively similar with cosine distance as the metric.
@@ 7 Related work
Implicit inductive biases towards simplicity: Highly over-parameterized deep networks sometimes generalize well even when they could memorize the entire dataset (Zhang et al., 2021). To make sense of these observations, a variety of works have proposed an implicit inductive bias for simplicity, either in the architectures themselves (Valle-Perez et al., 2018; Huh et al., 2021; Rahaman et al., 2019) or in combination with gradient-based learning dynamics (Tachet et al., 2018; Arora et al., 2019; Advani et al., 2020; Kalimeris et al., 2019; Lampinen and Ganguli, 2018; Shah et al., 2020; Lu et al., 2023; Xu et al., 2023). In particular, Valle-Perez et al. (2018) emphasize that networks with random parameters are biased towards computing simpler functions, which may contribute to the relative ease of learning simpler features—and the greater sparsity in representations of more complex features.
@@
Representational biases towards simple, common features: Several other empirical works have noted that representations can be biased towards capturing simpler features and suppressing irrelevant (or more complex) ones, including Hermann and Lampinen (2020) and more recent work on contrastive learning (Xue et al., 2023), and boolean feature learning (Qiu et al., 2024). In addition, several works have noted biases due to prevalence, e.g. that cosine similarity of word embeddings is biased by prevalence (Zhou et al., 2022), or that pruning selectively harms performance on rare exceptions within a category (Hooker et al., 2019). One earlier work on modularity (Csordás et al., 2021) noted in passing some biases in the proportion of network weights devoted to two different tasks. However, aside from that, most works have focused on the case where multiple features are used to predict the same label, and thus there is effective competition between features. Here we characterize these biases more systematically, in the case that all features are playing an important computational role. Moreover, we go beyond complexity to characterize other biases, such as the output position biases in transformers.
@@ 8 Discussion
For example, several recent works have proposed to tackle superposition in model activations via sparse dictionary learning (Bricken et al., 2023b; Cunningham et al., 2023). These methods use an autoencoder with a high-dimensional latent space and a sparsity penalty to process the internal representations, which can decompose them to yield more interpretable features. However, because the autoencoder is trained to reconstruct the internal representations via an $\ell^{2}$ loss, the objective favors reconstructing high-variance features; thus, these interpretability methods will likewise potentially be biased towards particular features and may neglect others that nevertheless play an important role in a model’s computations.
@@
It is natural to ask whether our terminology fits these levels of analysis appropriately. In our motivation, we described the different features as playing the same “computational” role, in the sense that the system is trained to output all of them. However, it would be reasonable to say that feature complexity, for example, changes the “computational” role that a feature plays in a model—it depends on where one draws the boundary between the computational and algorithmic. We have adopted a particular convention for clarity of presentation, but we do not think that adjusting this boundary and our terminology would fundamentally change our results or their implications.
@@
The work above is just one instance of a range of interventional methods that have been applied to analyze the computations of deep learning models (Geiger et al., 2021; Olsson et al., 2022; Geiger et al., 2024; Nanda et al., 2023; Conmy et al., 2023; Wu et al., 2024; Shah et al., 2024) and brains (Deisseroth, 2015; Liu et al., 2012). Again, by intervening on the system, these methods do yield stronger constraints, and allow stronger causal interpretations of which features are playing a particular role in a computation.
However, in practice, the hypotheses tested via these more rigorous interventional methods often originate from correlational analyses of the model’s representations (e.g. seeing that certain neurons are very responsive to a feature) or prior knowledge about a task (e.g. computational hypotheses about how the problem should be solved). Our experiments suggest that the first class of approaches may be biased in the kinds of causal hypotheses they generate, e.g. towards hypotheses about simpler features. Similarly, we may believe algorithmically that some features should be used to solve a task, and find representations that correlate with the feature and causally influence the models representations. However, in either case, it will be difficult to rule out the possibility that other subtle representational features are influencing the system’s computations in important ways. Furthermore, our analyses in Section 3.2 suggest that to fully identify the contribution of a feature that is nonlinearly involved in the network’s computations, we need to account for all the input patterns relevant to that feature—even with prior knowledge about which features are relevant to a task, this will be challenging in the general case. Thus, while causal methods are critical to thoroughly testing hypotheses about representation and computation, we do not believe they fully resolve the challenges raised here.
@@
Limitations & future directions: There are a number of limitations to our work that would be useful to address in the future. First, we focused on simple, synthetic datasets. While this choice allowed complete control over the tasks, the features, and their correlations, it is possible that biases would be different on more realistic datasets with larger sets of entangled features. (However, we note that concurrent work from Fel et al. (2024) finds compatible results for larger models trained on ImageNet.) Moreover, we focused on classification because, again, it made the computational solution more straightforward to analyze. However, it would be interesting to consider how these biases differ under other objectives—though an increasing amount of work has converged on some kind of (sequential) classification, e.g. language modeling. Moreover, while we did not see dramatic effects of parameter count in the regimes we considered, we did see larger biases in deeper models, and it is possible that qualitatively different biases would emerge in larger models; thus it would be useful to study these effects at scale in future work. Finally, the representations we consider are not richly contextual. However, transformers can learn to restructure their feature representations according to the task context (Li and McClelland, 2023); in future work it would be interesting to study how representation biases interact with more richly contextual computation.
@@ Acknowledgements
We thank Roma Patel, Aaditya Singh, Lukas Muttenthaler, and Ed Grefenstette for comments and suggestions on earlier drafts of the paper.
@@ B.3 MLP learning curves by training order
B.4 Feature complexity also biases sparsity of representations
Figure 22: Harder features also yield sparser representations. (Lines are means across 10 seeds, areas are 95%-CIs).
@@ B.4 Feature complexity also biases sparsity of representations
B.5 More features of varying complexity
@@
B.6 The effects of MLP architecture and other hyperparameters
@@
B.6.1 Dropout and optimizer interact for MLP architectures
@@
B.7 Transformer sequential learning curves
@@
B.8 Additional Dyck language results
@@
Figure 34: Test accuracy distributions (across runs) on superficial and structural features of sentences from Dyck-(20,10). (Bars are averages across 32 seeds, intervals are 95%-CIs.)
@@
B.9 Additional vision experiment results
@@
Figure 36: Learning curves for test accuracy and variance explained for CNNs and ResNets trained on the color-texture-shape tasks. (36(a)) For CNNs Color, which is learned first, quickly comes to dominate at the spatial pooling layer; the penultimate layer produces relatively less bias. (36(b)) For Resnets, color briefly dominates, but is quickly overtaken by shape. (Lines are averages across 32 seeds, areas are 95%-CIs.)
B.9.1 ResNet representational biases for shape are affected by the number of class labels and the presence of distractors
@@
Figure 37: Representational biases of ResNets trained to report binary classifications of the color, shape, and texture of a single object shift depending on the other objects in the image. (37(a)) with no other objects in the image, the representations still show shape bias, but more weakly than with the original 10-way classifications (Fig. 11(b)). (37(b)) With distractor objects in the image, the biases shift to favor color and texture. (Lines are averages across 32 seeds, areas are 95%-CIs.)
B.10 Keeping all but the top-$k$ principal components

### v2

@@ 5 Vision
We round out our experiments by considering a third large class of problems—computer vision—and some architectures typically used to solve them. Specifically, we construct synthetic datasets of images that show objects with distinct colors, shapes, and textures, and train ResNets (He et al., 2016) and other CNNs without residual connections to classify all of these features simultaneously. These experiments are again inspired by prior work finding that features like color are easier for models to learn, and will dominate when multiple features compete to predict a label (Hermann and Lampinen, 2020). Here, we investigate the models’ internal representations of these features when all features are learned.
We find that non-residual CNNs exhibit a strong bias towards the simplest feature (color, which can be decoded from any single pixel) at the final spatial pooling layer, and a somewhat weaker bias at the penultimate classifier layer (Fig. 11(a)). This is in keeping with the general pattern observed above, that models are biased towards representing simpler features more strongly, as well as with prior work. Intriguingly, however, ResNets exhibit the opposite bias—with representations that strongly favor shape (Fig. 11(b)). This is despite the fact that color remains the easiest feature, and shape the most difficult (slowest) feature to learn for both models (Fig. 38; as in Hermann and Lampinen 2020).
@@ 5.1 Relational vision tasks
Why do the basic feature biases of ResNets change direction in these tasks, compared to those above? In follow-up experiments (Appendix B.10.1) we show that switching to binary classification, and the addition of other objects to the image, may both contribute to reversing these biases. These results highlight the complex interactions that affect representational biases. However, the bias towards basic over relational features seems relatively more consistent.
@@ 6.1 The representations identified are causally sufficient and specific
The above analyses are correlational. However, most philosophical accounts (e.g. Cao, 2022; Baker et al., 2022) define representations partly in terms of their causal role in a computation. An increasing number of approaches in interpretability similarly use causal interventions to achieve more faithful interpretations of model internal representations (e.g. Geiger et al., 2021; Meng et al., 2022; Geiger et al., 2024). Here, we verify in one case that the representations identified in our analyses indeed play a causal role in the computations.
To do so, we take an MLP model (from the Fig. 2 experiments) at the end of training, manipulate its internal representations, and evaluate the causal effect on its output behavior. To manipulate model representations of a feature, we take the representation pattern predicted when that feature is present or absent and subtract those two patterns to create a present-absent difference vector. We then add (or subtract) a scalar multiple of this difference vector to the representation of each stimulus in the test dataset, and evaluate whether by doing so we can flip the label the model predicts for that feature on that stimulus.
@@ 6.2 Impact on model simplification for interpretability
Figure 14: If we keep only the top-$k$ principal components from a model’s representations, the easy feature will be preserved with a smaller $k$ than the hard feature.
@@
Though our experimental settings are somewhat different, our results hint at a possible basis for those phenomena. In particular, if the features required to solve the harder OOD splits are more complex, they may occupy relatively less variance, and therefore be more readily lost when keeping only the top principal components. We show a simple demonstration of this from one of our MLP models in Fig. 14: if we keep only two principal components from the models representations, the model can still identify the easy feature perfectly, while hard feature accuracy remains at chance until we keep at least four principal components. (In Appx. B.11 we show that dropping the top PCs yields the opposite pattern, with hard features proving more robust).
@@ 6.2.1 Impact on feature visualization
Given the above observations, clearly low-dimensional visualization may be biased towards certain features. We showed a simple example of this in Fig. 3, in which the top PCs are entirely clustered by the easy feature. In Appx. B.12 we elaborate these results, showing that the hard feature appears in the next PC. However, if models are trained with more and more easy features, the hard feature is pushed into higher and higher PCs. Visualizing with $t$-SNE yields somewhat better results with only a single easy feature—due to $t$-SNE’s nonlinear embedding and emphasis on neighbors—but once enough easy features are added it similarly becomes difficult to identify the hard feature in low-dimensional $t$-SNE plots as well, while the easy features show extremely clear clusters. Thus, feature visualizations may be inherently biased towards simpler features, those that are learned earlier, etc.
@@ 6.3 Impact on Representational Similarity Analysis
To illustrate this, in Fig. 15 we perform an RSA analysis between many pairs of MLP networks. Each MLP is trained to output two feature; we vary the complexity of these features across the different models. Ideally, the RSA would be more similar on the diagonal (for networks trained to perform the same two tasks) than off the diagonal, and those on the block diagonals (trained with just one feature that is the same) would be more similar than those off. However, the feature representation biases lead to a substantially more biased structure. For example, two models that are both computing SUM%2 over two sets of features will appear to be less similar to one another than each appears to be a network that is solely computing simple linear features. These results use a Euclidean metric for RSA, which is naturally sensitive to variance. However, in Appx. B.13 we show that the results are qualitatively similar with cosine distance as the metric.
@@ 7 Related work
Implicit inductive biases towards simplicity: Highly over-parameterized deep networks sometimes generalize well even when they could memorize the entire dataset (Zhang et al., 2021). To make sense of these observations, a variety of works have proposed an implicit inductive bias for simplicity, either in the architectures themselves (Valle-Perez et al., 2018; Huh et al., 2021; Rahaman et al., 2019) or in combination with gradient-based learning dynamics (Tachet et al., 2018; Arora et al., 2019; Advani et al., 2020; Kalimeris et al., 2019; Lampinen and Ganguli, 2018; Shah et al., 2020; Lu et al., 2023; Xu et al., 2024). In particular, Valle-Perez et al. (2018) emphasize that networks with random parameters are biased towards computing simpler functions, which may contribute to the relative ease of learning simpler features—and the greater sparsity in representations of more complex features.
@@
Representational biases towards simple, common features: Several other empirical works have noted that representations can be biased towards capturing simpler features and suppressing irrelevant (or more complex) ones, including Hermann and Lampinen (2020) and more recent work on contrastive learning (Xue et al., 2023), and boolean feature learning (Qiu et al., 2024). In addition, several works have noted biases due to prevalence, e.g. that cosine similarity of word embeddings is biased by prevalence (Zhou et al., 2022), or that pruning selectively harms performance on rare exceptions within a category (Hooker et al., 2019). A more recent study by Morwani et al. (2023) suggests that some representational biases may stem from networks learning maximum-margin solutions; likewise, Frei et al. (2023) theoretically study the behavior and representation learning of models trained on XOR tasks, with similar conclusions. One earlier work on modularity (Csordás et al., 2021) noted in passing some biases in the proportion of network weights devoted to two different tasks. However, aside from that, most works have focused on the case where multiple features are used to predict the same label, and thus there is effective competition between features. Here we characterize these biases more systematically, in the case that all features are playing an equally-important, independent computational role. Moreover, we go beyond complexity to characterize other biases, such as the output position biases in transformers.
@@ 8 Discussion
For example, several recent works have proposed to tackle superposition in model activations via sparse dictionary learning (Bricken et al., 2023b; Cunningham et al., 2023). These methods use a sparse autoencoder (SAE) with a high-dimensional latent space and a sparsity penalty to process the internal representations, which can decompose them to yield more interpretable features. However, because the autoencoder is trained to reconstruct the internal representations via an $\ell^{2}$ loss, the objective favors reconstructing high-variance features; thus, these interpretability methods will likewise potentially be biased towards particular features and may neglect others that nevertheless play an important role in a model’s computations.
@@
It is natural to ask whether our terminology fits these levels of analysis appropriately. In our motivation, we described the different features as playing the same “computational” role, in the sense that the system is trained to output all of them. However, it would be reasonable to say that feature complexity, for example, changes the “computational” role that a feature plays in a model—it depends on where one draws the boundary between the computational and algorithmic. We have adopted a particular convention for clarity, but we do not think that adjusting this boundary and our terminology would fundamentally change our results or their implications.
@@
The work above is one instance of many interventional methods that have been applied to analyze the computations of deep learning models (Geiger et al., 2021; Olsson et al., 2022; Geiger et al., 2024; Nanda et al., 2023; Conmy et al., 2023; Wu et al., 2024; Shah et al., 2024) and brains (Deisseroth, 2015; Liu et al., 2012). Again, by intervening, these methods do yield stronger constraints, and allow stronger causal interpretations of which features are playing a particular role in a computation.
However, in practice, the hypotheses tested via these more rigorous interventional methods often originate from correlational analyses of representations (e.g. seeing that certain neurons are highly responsive to a feature) or prior knowledge about a task (e.g. computational hypotheses about how the problem should be solved). Our experiments suggest that the first class of approaches may be biased in the kinds of causal hypotheses they generate, e.g. towards hypotheses about simpler features. Similarly, we may believe algorithmically that some features should be used to solve a task, and find representations that correlate with the feature and causally influence the models representations. However, in either case, it will be difficult to rule out the possibility that other subtle representational features are influencing the system’s computations in important ways. Furthermore, our analyses in Section 3.2 suggest that to fully identify the contribution of a feature that is nonlinearly involved in the network’s computations, we need to account for all the ways that feature could be represented—even with prior knowledge about which features are relevant to a task, this will be challenging in the general case. Thus, while causal methods are critical to thoroughly testing hypotheses about representation and computation, we do not believe they fully resolve the challenges raised here.
@@
Limitations & future directions: There are a number of limitations to our work that would be useful to address in the future. First, we focused on simple, synthetic datasets. While this choice allowed complete control over the tasks, the features, and their correlations, it is possible that biases would be different on more realistic datasets with larger sets of entangled features. (However, concurrent work from Fel et al. (2024) finds compatible results for larger models trained on ImageNet.) Moreover, we focused on classification because, again, it made the computational solution more straightforward to analyze. However, it would be interesting to consider how these biases differ under other objectives—though an increasing amount of work has converged on some kind of (sequential, soft) classification, e.g. language modeling. Moreover, while we did not see dramatic effects of parameter count in the regimes we considered, we did see larger biases in deeper models, and it is possible that qualitatively different biases would emerge in larger models; thus it would be useful to study these effects at scale in future work. Finally, the representations we consider are not richly contextual. However, transformers can learn to restructure feature representations according to the task context (Li and McClelland, 2023); in future work it would be interesting to study how representation biases interact with more richly contextual computation.
@@ Acknowledgements
We thank Roma Patel, Aaditya Singh, Spencer Frei, Dan Friedman, Lukas Muttenthaler, Asma Ghandeharioun, and Ed Grefenstette for comments and suggestions on earlier drafts of the paper.
@@ B.3 MLP learning curves by training order
B.4 Per-unit representation variance analysis
In this section we present analyses of how each of the 64 units in the penultimate layer of the model represents each feature for a single run of the basic model trained on the linear and sum mod 2 tasks. In Fig. 22 we show snapshots of the variance explained by each feature for each unit, at various salient stages in training. In Fig. 23 we show individual learning curves for all the units. The qualitative patterns are similar to the aggregate case, but there are interesting individual effects to observe. Most notably, the units that end up carrying the most variance about the difficult feature are also those that initially carry the most variance about the easy feature, and the variance they carry about the easy feature is inflated as the hard feature is learned. Thus, most of the individual units also continue to carry more variance due to the easy feature. Only a subset of the initially-easy-selective neurons acquire some hard-relevant signal later (which likely contributes to the sparsity difference observed below). However, a few neurons that initially carried at least as much signal about the hard inputs as easy ones do tend to become somewhat more selective for the hard input features as the harder task is learned. Furthermore, many of the units carry non-trivial signal about other inputs (or nonlinear patterns of the feature-relevant inputs) that is also amplified as the other features are learned — either because those inputs are weakly correlated with a task feature within the dataset by chance, or just due to the learning dynamics.
Figure 22: The distribution of variance explained by each feature across the 64 units of an MLP at various stages of training. (The ordering of units is consistent across the plots, and is determined by the final plot.)
Figure 23: Raw representation variance plotted across learning for each of the 64 units in the penultimate layer of an MLP. Note that this figure plots raw variance explained, without normalizing by the total variance of the unit, to facilitate comparing across units. Note also that the vertical axis is log-scaled; where the variance explained on the test set was $\leq 0$ it is omitted from the plot.
B.5 Feature complexity also biases sparsity of representations
Figure 24: Harder features also yield sparser representations. (Lines are means across 10 seeds, errorbars are 95%-CIs).
@@ B.5 Feature complexity also biases sparsity of representations
B.6 More features of varying complexity
@@
B.7 The effects of MLP architecture and other hyperparameters
@@
B.7.1 Dropout and optimizer interact for MLP architectures
@@
B.8 Transformer sequential learning curves
@@
B.9 Additional Dyck language results
@@
Figure 36: Test accuracy distributions (across runs) on superficial and structural features of sentences from Dyck-(20,10).
@@
B.10 Additional vision experiment results
@@
Figure 38: Learning curves for test accuracy and variance explained for CNNs and ResNets trained on the color-texture-shape tasks. (38(a)) For CNNs Color, which is learned first, quickly comes to dominate at the spatial pooling layer; the penultimate layer produces relatively less bias. (38(b)) For Resnets, color briefly dominates, but is quickly overtaken by shape. (Lines are averages across 32 seeds, areas are 95%-CIs.)
B.10.1 ResNet representational biases for shape are affected by the number of class labels and the presence of distractors
@@
Figure 39: Representational biases of ResNets trained to report binary classifications of the color, shape, and texture of a single object shift depending on the other objects in the image. (39(a)) with no other objects in the image, the representations still show shape bias, but more weakly than with the original 10-way classifications (Fig. 11(b)). (39(b)) With distractor objects in the image, the biases shift to favor color and texture. (Lines are averages across 32 seeds, areas are 95%-CIs.)
B.11 Keeping all but the top-$k$ principal components